3. Measurements show that 500 muon fall on the top of Mt. Washington altitude 2000m. Muons travel at 99% the speed of light or 0.99 x 3.0 x 108 m/s.

a. Calculate the time in microseconds it would take muon to travel from the top of mt. Washington to its base.

b. Use your calculation and the muon clock graph to find how many muons should reach the bottom of Mt. Washington.

4. Experiments show that the actual number of muons that reach the base of Mt. Washington is 400.

a. According to your muon clock graph, how much time has elapsed if 400 muons reach the base of Mt. Washington?

b. By what factor do the times you found differ?

c. Suggest an explanation for this difference.



1
Expert's answer
2022-03-15T16:49:44-0400

"t=\frac sv=6.73~\mu s,"

"N=N_0(1-2^{-\frac tT})=475."